OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)? A.2 B.-4/5 C.5/4 D.-2

OpenStudy (anonymous):

@Angel_Bear19

OpenStudy (eric_d):

same method

OpenStudy (eric_d):

Y2-Y1/X2-X1

OpenStudy (eric_d):

=m

OpenStudy (anonymous):

1- (-5) ------ 4-1

OpenStudy (anonymous):

What do you think the answer is Bree?

OpenStudy (anonymous):

-5 -- 4

OpenStudy (eric_d):

nope

OpenStudy (anonymous):

when its a negative minus a negative you add, try again

OpenStudy (eric_d):

(-)(-)=(+) @Bree123

OpenStudy (anonymous):

6 -- 3

OpenStudy (eric_d):

yes

OpenStudy (anonymous):

Right! which simplifies to 2 (:

OpenStudy (eric_d):

simplify it

OpenStudy (anonymous):

2

OpenStudy (eric_d):

yup

OpenStudy (anonymous):

slope(m)= (Y2-Y1)/ (X2-X1) M= (1-(-5))/(4-1) m= (1+5)/(3) m= 6/3 m=2